One side of a triangle is half the longest side. The third side is 10 meters less than the longest side. The peremiter is 53 feet. Find the shortest side.

Answer by mananth(16946)

let longest side = x
one side = x/2
third side =x-10
...
x+x/2 + x-10=53
+10
2x+x/2 +10-10=53+10
2x+x/2 = 63
(4x+x)/2 =63
LCD =2
5x/2 = 63
multiply by 2
5x= 2*63
/5
x= 2*63/5
=25.2
..
x/2 = 25.2 /2 = 12.6
x-10 = 25.2 -10=15.2
The sides are
25.2 , 12.6, 15.2
Check
Their sum = 53